George Rentals offers machine rental services for concrete cutting. Consider the following costs of the company over the relevant range of 4,000 to 10,000 hours of operating time for its concrete cutting equipment. Hours of Operating Time 4,000 5,000 8,000 10,000 Total Costs: Variable Costs $ 20,000 ? ? ? Fixed Costs 82,000 ? ? ? Total Costs $102,000 ? ? ? Cost per hour: Variable cost ? ? ? ? Fixed cost ? ? ? ? Total cost per hour ? ? ? ? What is the estimated total cost per hour at a volume of 8,000 hours?
Answer #1
Variable cost per hour 5.00 =20000/4000
Estimated variable cost at 8000 hours 40000 =8000*5
Estimated fixed cost at 8000 hours 82000
Estimated total cost at 8000 hours 122000
Estimated total cost per hour at a volume of 8,000 hours 15.25 =122000/8000